Solve (log (base 5) x)^2 =( log (base 5) x) + 3
Answer by stanbon(75887)

Solve (log (base 5) x)^2 =( log (base 5) x) + 3
-----------
log5(x^2) = log5(x) + 3
------
log5(x^2)-log5(x) = 3
----
log5(x^2/x) = 3
----
log5(x) = 3
-----
x = 5^3
-----
x = 125
==============
Cheers,
Stan H.
==============